What is the answer to 4/20 simplified as a fraction?
1
Expert's answer
2020-02-26T10:21:20-0500

Solution.

"\frac{4}{20}=\frac{4\times1}{4\times5}=\frac{1}{5}"

Answer: "\frac{1}{5}."